# AI-driven inverse design of materials: Past, present and future

### v1

@@ Abstract
The discovery of advanced materials is the cornerstone of human technological development and progress. The structures of materials and their corresponding properties are essentially the result of a complex interplay of multiple degrees of freedom such as lattice, charge, spin, symmetry, and topology. This poses significant challenges for the inverse design methods of materials. Humans have long explored new materials through a large number of experiments and proposed corresponding theoretical systems to predict new material properties and structures. With the improvement of computational power, researchers have gradually developed various electronic structure calculation methods, particularly such as the one based density functional theory, as well as high-throughput computational methods. Recently, the rapid development of artificial intelligence technology in the field of computer science has enabled the effective characterization of the implicit association between material properties and structures, thus opening up an efficient paradigm for the inverse design of functional materials. A significant progress has been made in inverse design of materials based on generative and discriminative models, attracting widespread attention from researchers. Considering this rapid technological progress, in this survey, we look back on the latest advancements in AI-driven inverse design of materials by introducing the background, key findings, and mainstream technological development routes. In addition, we summarize the remaining issues for future directions. This survey provides the latest overview of AI-driven inverse design of materials, which can serve as a useful resource for researchers.
@@ 1 Introduction
Figure 2: Materials Science Research Paradigms This figure illustrates the evolution of research paradigms in materials science, emphasizing key milestones across various approaches. It highlights the increasing role of artificial intelligence in driving future materials discovery, with AI becoming the dominant force in shaping the field. The experiment-driven paradigm is exemplified by Marie Curie’s Nobel Prize-winning discovery of radium and polonium. The theory-driven paradigm is represented by Dirac’s equation, which predicted the existence of the positron, as well as the quantum spin Hall effect. The computation-driven paradigm is demonstrated through DFT calculations applied to materials such as MgB2, hydride superconductors, and graphene. Finally, the AI-driven paradigm showcases recent breakthroughs in artificial intelligence, including AlphaFold2, GPT-4, and AI-accelerated materials discovery, signaling the frontier of research in the field.
@@ 2.2 Magnetic Materials
Based on magnetic properties, magnetic materials can be categorized three types: ferromagnetic [88], antiferromagnetic [89], and altermagnetic [90, 91, 92, 93, 94, 95, 96, 97]. The difference between ferromagnetic materials and antiferromagnetic materials lies in the arrangement of the magnetic moments within a material and the resulting magnetic behavior. Ferromagnetic materials exhibit a wide range of magnetic domains [98], where the magnetic moments within each domain are aligned in the same direction. However, the magnetic moments in different domains may be oriented differently. Under the influence of an external magnetic field, ferromagnetic materials can be strongly magnetized, and they possess a high coercivity, meaning a strong reverse magnetic field is required to change the direction of magnetization. After the removal of the external magnetic field, ferromagnetic materials maintain a certain level of remanence and exhibit high permeability. The altermagnetic materials, despite exhibiting antiferromagnetic compensated magnetic order, show macroscopic time-reversal symmetry-breaking phenomena and spin polarization similar to ferromagnet. This seemingly contradictory behavior has led scientists to reclassify these materials based on spin-symmetry principles, thereby defining altermagnetism as a new magnetic phase. As research on altermagnetic materials deepens, more material candidates have been discovered with these properties, including insulators, semiconductors, metals, and high-temperature superconductors [91].
@@ 2.5 Two-dimensional Materials
Due to the extremely thin thickness, the movement of electrons in 2D materials is highly constrained, leading to the quantum confinement effect. This effect often results in novel physical properties, making 2D materials of great interest in fields such as electronics [153], spintronics [154], valleytronics [155], optoelectronics [156], twistronics [157], and slidetronics [158]. Common examples of 2D materials include graphene, which consists of a single layer of carbon atoms arranged in a honeycomb lattice [159]. Another important class of 2D materials is transition metal dichalcogenides, typically composed of one layer of transition metal atoms (such as $Mo$ or $W$) sandwiched between two layers of chalcogenide atoms (such as $S$, $Se$ or $Te$[160]. For van der Waals heterostructures, interlayer coupling can be performed by regulating the interaction between different two-dimensional material layers [161].By changing the number of layers and the stacking order, the properties of the van der Waals heterostructures can be fine-tuned, and the 2D materials can be turned into topological insulators through specific layering structures [162].
@@ 3 The Development of AI Methods in Materials Science
Figure 4: The rapid advancement of AI technologies accelerates materials discovery in various ways. Orange indicates traditional machine learning; green represents the development of invariant GNNs and their application in discriminative AI for predicting material properties; blue denotes the progress of equivariant GNNs, generative AI techniques, and their integration for structural predictions. Gray signifies the role of LLMs in expediting materials discovery.
@@ 3.3 Discriminative AI
CDVAE. The crystal diffusion variational autoencoder (CDVAE) [351] stands as a prominent early algorithm developed to generate stable periodic structures of materials. It employs a combination of joint equivariant diffusion and SE(3)-equivariant graph neural networks, explicitly encoding interactions across periodic boundaries while preserving permutation, translation, rotation, and periodic invariance. Furthermore, CDVAE introduces a noise conditional score network (NCSN) [346] as the VAE decoder to generate more realistic material structures, utilizing Langevin dynamics to refine the generation process. The algorithm also establishes three benchmark datasets specifically for material generation, along with a set of physically meaningful evaluation tasks and metrics. Extensive experiments validate superior performance of CDVAE in tasks such as material reconstruction, generation, and property optimization, underscoring its significant potential in advancing the field of materials discovery. The workflow of CDVAE consists of several key steps. First, a periodic material encoder is employed to map the material ${\bm{M}}$ into a latent representation. Then, three independent MLPs predict the composition ${\bm{c}}$, lattice ${\bm{L}}$, and number of atoms $N$, respectively. A conditional score matching decoder with equivariance is used to denoise the atomic coordinates and the probability distribution of atom types. To account for the periodicity of atomic coordinates, a constraint is incorporated into the loss function. Finally, during each denoising step, atomic types and coordinates are updated using Langevin dynamics.
@@ 5 Conclusion
Survey Writing. This survey was planned during a discussion meeting held by our research team, with the objective of summarizing the latest advancements in AI-driven inverse design of materials into a highly readable report. This is the first draft of our survey, completed on November 13, 2024, and it will be updated in a timely manner as needed.

### v2

@@ Abstract
The discovery of advanced materials is the cornerstone of human technological development and progress. The structures of materials and their corresponding properties are essentially the result of a complex interplay of multiple degrees of freedom such as lattice, charge, spin, symmetry, and topology. This poses significant challenges for the inverse design methods of materials. Humans have long explored new materials through a large number of experiments and proposed corresponding theoretical systems to predict new material properties and structures. With the improvement of computational power, researchers have gradually developed various electronic structure calculation methods, such as the density functional theory and high-throughput computational methods. Recently, the rapid development of artificial intelligence technology in the field of computer science has enabled the effective characterization of the implicit association between material properties and structures, thus opening up an efficient paradigm for the inverse design of functional materials. A significant progress has been made in inverse design of materials based on generative and discriminative models, attracting widespread attention from researchers. Considering this rapid technological progress, in this survey, we look back on the latest advancements in AI-driven inverse design of materials by introducing the background, key findings, and mainstream technological development routes. In addition, we summarize the remaining issues for future directions. This survey provides the latest overview of AI-driven inverse design of materials, which can serve as a useful resource for researchers.
@@ 1 Introduction
Figure 2: Materials Science Research Paradigms This figure illustrates the evolution of research paradigms in materials science, emphasizing key milestones across various approaches. It highlights the increasing role of artificial intelligence in driving future materials discovery, with AI becoming the dominant force in shaping the field. The experiment-driven paradigm is exemplified by Marie Curie’s Nobel Prize-winning discovery of radium and polonium. The theory-driven paradigm is represented by Dirac’s equation, which predicted the existence of the positron and the quantum spin Hall effect. The computation-driven paradigm is demonstrated through DFT calculations applied to materials such as MgB2, hydride superconductors, and graphene. Finally, the AI-driven paradigm showcases recent breakthroughs in artificial intelligence, including AlphaFold2, GPT-4, and AI-accelerated materials discovery, signaling the frontier of research in the field.
@@
With the development of technology, a new research field has emerged in materials science known as the inverse design of materials. Unlike traditional trial-and-error methods, it utilizes complex computational techniques to design materials with specific properties from the ground up. The process involves several key steps: first, defining the target properties or functionalities; then, selecting an appropriate design space; followed by using modeling and simulation tools (such as DFT and finite element analysis) to simulate material properties; optimizing the material structure through algorithms to minimize the difference between simulated properties and target goals; finally, validating the design through experimental verification, followed by iterative refinement based on experimental results. This approach is widely applied in areas such as optics, electronics, energy storage, catalysis, and composite materials, significantly accelerating the development of new materials. However, inverse design also faces challenges such as computational cost, data quality, and experimental validation. With the advancement of computational power and optimization algorithms, materials inverse design is expected to play an increasingly important role in the future development of materials science.
@@ 2.2 Magnetic Materials
Based on magnetic properties, magnetic materials can be categorized three types: ferromagnetic [88], antiferromagnetic [89], and altermagnetic [90, 91, 92, 93, 94, 95, 96, 97, 98]. The difference between ferromagnetic materials and antiferromagnetic materials lies in the arrangement of the magnetic moments within a material and the resulting magnetic behavior. Ferromagnetic materials exhibit a wide range of magnetic domains [99], where the magnetic moments within each domain are aligned in the same direction. However, the magnetic moments in different domains may be oriented differently. Under the influence of an external magnetic field, ferromagnetic materials can be strongly magnetized, and they possess a high coercivity, meaning a strong reverse magnetic field is required to change the direction of magnetization. After the removal of the external magnetic field, ferromagnetic materials maintain a certain level of remanence and exhibit high permeability. The altermagnetic materials, despite exhibiting antiferromagnetic compensated magnetic order, show macroscopic time-reversal symmetry-breaking phenomena and spin polarization similar to ferromagnet. This seemingly contradictory behavior has led scientists to reclassify these materials based on spin-symmetry principles, thereby defining altermagnetism as a new magnetic phase. As research on altermagnetic materials deepens, more material candidates have been discovered with these properties, including insulators, semiconductors, metals, and high-temperature superconductors [91].
@@ 2.5 Two-dimensional Materials
Due to the extremely thin thickness, the movement of electrons in 2D materials is highly constrained, leading to the quantum confinement effect. This effect often results in novel physical properties, making 2D materials of great interest in fields such as electronics [154, 155, 156], spintronics [157], valleytronics [158], optoelectronics [159], twistronics [160], and slidetronics [161]. Common examples of 2D materials include graphene, which consists of a single layer of carbon atoms arranged in a honeycomb lattice [162]. Another important class of 2D materials is transition metal dichalcogenides, typically composed of one layer of transition metal atoms (such as Mo or W) sandwiched between two layers of chalcogenide atoms (such as S, Se or Te) [163]. For van der Waals heterostructures, interlayer coupling can be performed by regulating the interaction between different two-dimensional material layers [164].By changing the number of layers and the stacking order, the properties of the van der Waals heterostructures can be fine-tuned, and the 2D materials can be turned into topological insulators through specific layering structures [165].
@@ 3 The Development of AI Methods in Materials Science
Figure 4: The rapid advancement of AI technologies accelerates materials discovery in various ways. Blue indicates traditional machine learning; green represents the development of invariant GNNs and their application in discriminative AI for predicting material properties; deep blue denotes the progress of equivariant GNNs, generative AI techniques, and their integration for structural predictions. Orange signifies the role of LLMs in expediting materials discovery.
@@ 3.3 Discriminative AI
[…]
@@
Figure 5: The general process of materials inverse design, which is divided into two main parts: the AI-driven theoretical calculation part and the experimental validation part. The theoretical calculation part is further subdivided into material design and generation, high-throughput screening, and computational modeling.
InvDesFLow. The InvDesFLow [72] develop an AI-driven workflow for discovering high-$T_{c}$ superconductors that are not present in any existing database. In this survey, we summarized the general process for inverse design of materials, as shown in Figure 5. The process is divided into two main parts: theoretical calculations and experimental validation. The theoretical calculations are further subdivided into material design and generation, high-throughput screening, and computational models. Unlike traditional reverse design, the entire theoretical part is fully AI-driven, with AI implementing the entire theoretical simulation process.
@@ 5 Conclusion
Survey Writing. This survey was planned during a discussion meeting held by our research team, with the objective of summarizing the latest advancements in AI-driven inverse design of materials into a highly readable report. Then, we have extensively revised the writing and contents in several passes. Due to the space limit, we can only include a fraction of existing AI methods in Figure 4 by setting the selection criterion. We release the initial version on November 14, 2024, and this latest version on November 26, 2024.
@@
Update log. In this part, we regularly maintain an update log for the submissions of this survey to arXiv:
$\bullet$ First release on November 14, 2024: the initial version.
$\bullet$ Update on November 26, 2024: add several related work in Section 2 and Section 4, revise Figure 4 and Table II, add the Figure 5, improve the writing, and correct some minor errors.